OpenStudy (anonymous):

If no digit may be used more than once, how many 2-digit numbers can be formed using only the digits 5, 6, 3, and 2?

OpenStudy (raden):

use 4P2

OpenStudy (raden):

= 4!/(4-2)! = 4!/2! = ...

OpenStudy (anonymous):

Thanks, I feel a little stupid for not seeing the answer to that one in the first place.

OpenStudy (raden):

no problem... 4! means 4 * 3 * 2 * 1 = 24 2! = 2 * 1 = 2 so, 4!/2! = 24/2 = 12